OpenStudy (anonymous):

medal rewarded! solve each equation of x. Use calculators approximate to 2 dec. places a. 8x^9=6x^6 b.15x-3=10x^-2 c.200x-1=125x-3

OpenStudy (anonymous):

c. is 200x^-1=125x^-3 typed that in wrong

OpenStudy (anonymous):

what do you mean by ^?

OpenStudy (anonymous):

to the power of

OpenStudy (anonymous):

here ^ means "to the power of" a) x^3 = 6/8 x = cuberoot(6/8)

OpenStudy (anonymous):

i'm really confused :/